OpenStudy (anonymous):

The equation y = 16x shows that Carla earns $16 for each hour she works. Which table represents the equation y = 16x?

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

y = 16x y = 16(0) y = 0 (0, 0) y = 16(1) y = 16 (1, 16) y = 16(2) y = 32 (2, 32) y = 16(3) y = 48 (3, 48)

OpenStudy (anonymous):

ty xD